For the equation CH4 + 2O2 →CO2 + H2O + 803.1 kJ, how much heat will be released if 2.0 x 10^1 g of CH4 react with sufficient O2? _______


A. 883 kJ

B. 883 kJ/mol

C. 1.0 x 10^3 kJ

D. 1.0 x 10^3 kJ/mol

E. None of the Above


1
Expert's answer
2021-05-13T02:06:23-0400

n(CH4) = m/M = 20g/16g/mol = 1.25 mol;


∆H' = 1.25×∆H = 1.25×803.1 = 1,004 kJ.


Answer: C. 1,004 kJ